Find the maclaurin series for this equation, written in sigma notation:

x^3/(1-3x^2)
Answer #1

Solution: 1 f(x)=1-32 (1-3) 1-3x2 DO 1 We have power series for = 1+x+x2+x+x* +... = x* for Ixl<1 1-X For g(x)= 7+34 1 1 1-x